### v3

# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

Abstract
Recent advances in large Language Models (LLMs) have revolutionized mobile robots, including unmanned aerial vehicles (UAVs), enabling their intelligent operation within Internet of Things (IoT) ecosystems. However, LLMs still face challenges from logical reasoning and complex decision-making, leading to concerns about the reliability of LLM-driven UAV operations in IoT applications. In this paper, we propose a LLM-driven closed-loop control framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator. Our framework transforms numerical state observations from UAV operations into natural language trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation. Our framework also enables a simulation-based refinement process, and hence eliminates the risks to physical UAVs caused by incorrect code execution during the refinement. Extensive experiments on UAV control tasks with different complexities are conducted. The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline approaches in terms of success rate and completeness with the increase of task complexity.
Index Terms:
UAV, Large Language Model, Robotics, IoT.
I Introduction
Uavs, or drones, are increasingly being integrated into IoT ecosystems to support various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Designing UAVs that have intelligence to perform complex tasks in these applications correctly remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [2, 3, 4, 5] have demonstrated remarkable proficiency in robotic areas such as planning [6], navigation [7, 8], communication [9], and automation [10]. With LLMs’ exceptional natural language (NL) understanding and context generation capabilities, LLMs offer a promising avenue for robots with greater intelligence. Motivated by these developments, research efforts have been made towards the reliable integration of LLMs and UAVs [11, 12, 13, 14, 15], to enable LLM-driven UAV operations.
Despite LLMs’ potential for various robotic tasks, LLMs continue to face challenges from logical reasoning [16], undermining their reliability in tasks that require complex decision-making. Crucially, unlike text generation applications where semantic-level accuracy is sufficient, UAV operation requires the precise execution of sequential actions. Erroneous or invalid control code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes. Recent studies leverage prompt engineering [17] to enhance LLM in translating NL instructions into control code [11, 18, 15]. However, these approaches adopt open-loop designs and lack mechanisms to verify the correctness of the synthesized code, leaving questions such as cumulative error and temporal action consistency unsolved [19].
Figure 1: Illustration of LLM-driven closed-loop UAV Operation and Refinement.
Inspired by control theory, more recent studies have adopted a closed-loop architecture that enables evaluation of generated code and provides feedback for refinement, thereby steering robot execution towards prescribed task objectives [20, 21, 22, 23, 24, 25, 26]. Nevertheless, challenges are prone. First, existing approaches prompt LLM with numerical state observations in scenarios that require insights into robot dynamics for evaluation [24, 26]. Given that LLMs exhibit limited numerical reasoning and processing capabilities [27], an LLM-driven evaluation may overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of the robot’s movements and dynamics, leading to unreliable code refinement feedback. Second, current closed-loop approaches require physical robot execution to generate state observations for evaluation [22, 24]. Such an approach may be acceptable for ground robots where occasional errors rarely cause irreparable damage. However, it poses risks for aerial platforms, as even slight control deviations can precipitate UAV crashes.
To address these challenges, we propose a novel closed-loop feedback and refinement framework that leverages semantic observations to enhance the reliability of LLM-driven UAV control as depicted in Fig. 1. Our framework comprises two LLM modules: a Code Generator, which synthesizes or refines UAV operation code based on task description or evaluation feedback; and an Evaluator, which assesses code performance and provides feedback to guide code refinement. The process unfolds as follows: 1) The code generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into an NL trajectory observation; 3) The evaluator compares NL trajectory observation against the original task specification and returns evaluation feedback; 4) If the evaluation confirms the trajectory meets the task objectives, the code will deploy on the UAV for mission execution. Otherwise, the evaluation feedback is returned to the code generator for iterative refinement. By introducing NL trajectory representations rather than numerical state information, our method significantly enhances the evaluator’s comprehension of the UAV’s state dynamics, thereby improving the reliability and precision of the evaluation feedback. Moreover, our method enables the refinement process in simulation, which eliminates the risks to UAVs caused by incorrect code execution during refinement, and hence supports safe and efficient development.
We conduct extensive experiments on UAV operation tasks with varying levels of complexity to evaluate the effectiveness of our proposed framework. The experimental results show that our framework achieves a $100\%$ success rate and $100\%$ completeness for basic operations. For complex tasks, our framework achieves an $85\%$ success rate and $98.5\%$ completeness. Compared with baseline approaches [15, 25, 24], our framework greatly improves both success rate and completeness when the task complexity increases. The contributions of our work can be summarized as follows:
  • Closed-Loop Refinement: A novel closed-loop feedback and refinement framework with a code generator and an evaluator designed to enhance the reliability of LLM-driven UAV control.
  • Semantic Encoding: A novel algorithm for converting numerical state observations into step-wise NL trajectory descriptions, thereby enhancing the evaluative reasoning process.
  • Safe Development: A simulation-based LLM-generated UAV code evaluation that replaces physical deployment to prevent potential UAV crashes in code refinement.
  • Comprehensive Evaluation: A comprehensive experimental evaluation that demonstrates the performance of our framework with a detailed analysis of design factors.
The remainder of this paper is structured as follows: Section II reviews and discusses existing research. Section III presents our problem formulation, which is followed by the detailed design of our proposed framework in Section IV. Section IV-D provides the evaluation and analysis of our framework. This paper is concluded in Section VI.
II Related Work
II-A LLM for Robotics
LLMs have demonstrated exceptional capabilities in semantic understanding and context generation, making them invaluable in robotics. ChatGPT for Robotics [11] and Code as Policies [18] have shown significant progress in LLM to control various types of robots. In the domain of aerial platforms, TypeFly [12] introduces an end-to-end LLM framework for piloting drones via human languages, while REAL [28] integrates LLMs into a mission planning and control framework. Moreover, continual advancements in LLMs’ reasoning and code generation capabilities [5, 29] are opening new opportunities for a variety of LLM-driven robotic tasks. Despite these advances, robotic performance continues to be limited by the reasoning capacity of LLMs, which frequently demonstrate unreliability when faced with tasks requiring complex reasoning [11].
II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [17] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [15]. Prompt engineering typically leverages two core techniques: in-context learning [30] and Chain-of-Thought (CoT) reasoning [31]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [18, 32]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [21, 32]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation in LLM-driven UAV control.
II-C LLM-based Feedback, Evaluation, and Refinement
ChatGPT for Robotics emphasizes the necessity of human oversight for code generation [11], reflecting the fact that robotic performance remains constrained by LLMs’ reasoning limitations. Moreover, LLMs exhibit sensitivity to input perturbations [33, 34, 35]. Thus, recent research efforts adopt closed-loop refinement strategies to verify the LLM generation for improving the reliability of LLM-driven control. For instance, PluginSafety [36] designs a module that enforces predefined logical constraints evaluation to ensure LLM-generated content adheres to predefined constraints. More recent studies employ corrective feedback to iteratively refine generated code. Some integrate human or automated checker feedback [20, 21, 37, 38] to re-prompt the LLM and improve reasoning accuracy. While others monitor the robot status during execution and provide feedback, such as runtime errors, robot constraints violations [39, 22, 23, 24]. The evaluation processes generally fall into two categories: self-evaluation and external evaluation. SELF-REFINE [25, 39] leverages a single LLM to assess and refine its own outputs based on feedback, whereas external evaluation approaches employ a secondary LLM or external evaluator to generate corrective feedback [36, 22, 23, 24]. In our approach, numerical state information is first transformed into NL observations, and then we use an external LLM-evaluator to provide feedback for iterative refinement.
Figure 2: An example of how our LLM-driven closed-loop feedback and refinement framework works during solving a task that requires examining a square area. In the first loop, the evaluator identifies deviations in the NL trajectory observation (action 12), providing error feedback into the code generator to refine the code. In the second loop, the evaluator confirms refined code matches the task, and the code is deployed on the UAV for mission execution.
III Problem Formulation
Given a UAV operation task description, the framework proposed in this paper aims to use LLMs to reason through the task description and generate accurate operation code after observation, evaluation, feedback, and refinement. Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [15] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state. For a current state $S$, an action $a\in A$ is selected from the skill set $A$, leading to a deterministic state transition: $T:S\times A\rightarrow S$ that determines the next state. A candidate solution to problem $P$ is a sequence of actions $l=(a_{1},a_{2},\dots,a_{n})$ that control the UAV from the initial state $s_{0}$ through intermediate states until reaching a final state $s_{n}$. The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback. If $F$ is ”YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
IV Methodology
IV-A Overview
The overall idea of our proposed reliable LLM-driven closed-loop control framework for UAV operations is illustrated in Fig. 2 and Algorithm 1. Given a UAV operation task description by the user, the LLM configured as the code generator with enhanced reasoning capabilities towards UAV operation first produces the initial version of the operation code. A simulation-based execution of the code is then performed to obtain the numerical state observations. To close the gap between numerical state observations and the LLM-driven evaluation, a state transformation is designed that transforms numerical state observations into an NL trajectory observation. The evaluator analyzes the trajectory observation together with the task description. If the trajectory satisfies all task objectives in evaluation, the operation code is ready for deployment; otherwise, the evaluator provides feedback to the code generator for refinement. This process iterates until the task objectives are achieved or the maximum number of iterations is reached. We now provide the detailed design of our code generation, trajectory observation, evaluation, feedback, and refinement in the following sections.
Input: $Task$: Task description
Output: $Code$: Code script
$Code\leftarrow$ Code Generator$(Task)$
$itr\leftarrow 1$
while $itr\leq$ maximum round do
       $O\leftarrow$ State Transformation$(Code)$
      $F\leftarrow$ Evaluator$(O,Task)$
      if “YES” in $F$ then
             break
      else
             $Code\leftarrow$ Code Generator$(F)$
       end if
      $itr\leftarrow$ $itr+1$
end while
return $Code$
Algorithm 1 Closed-loop Feedback and Refinement
IV-B Code Generation
The LLM-driven code generation in our framework is achieved by configuring a general-purpose LLM agent (code generator) with UAV operation-related system prompts. We adopt prompt engineering strategies proposed by the GSCE framework [15], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation. Specifically, the system prompts comprise guidelines, skill APIs, constraints, and examples. The guidelines shape the LLM’s role for code generation, the skill APIs empower LLM to employ appropriate functions, the constraints regulate LLM with robotic policies, and the examples are pairs of query-response that serve as illustrations. These strategies serve as the basic building blocks to support the LLM-driven UAV operation code generation. On top of that, our code generation also utilizes constructive feedback from our evaluation design in section IV-D. The feedback helps the code generator obtain a more comprehensive understanding of the dynamics and interconnections among UAV operations implied by the task description, which enhances the LLM’s reasoning for code generation and code refinement.
IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [24, 22] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of numerical state tuples [27].
To bridge this gap, we transform each numerical state observation into a step-wise NL description and aggregate these into an NL trajectory observation. As shown in Fig. 1, NL description captures the semantic meaning of the state changes, such as “Move 5 meters West while facing North, the UAV moves to [0, 5]”. This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for code refinement.
Given the LLM-generated UAV operation code executed in simulation, Algorithm 2 presents the transformation from numerical state observations to the NL trajectory observation $O$. For each action identified in the $Code$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations. Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the difference between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $\Delta s=s_{\mathrm{current}}-s_{\mathrm{last}}$. For each element in the state difference $\Delta s$, the algorithm appends the corresponding NL description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$. The algorithm also appends the UAV’s NED position to $O$. Upon completion of all actions, the algorithm returns the complete NL trajectory observation $O$.
Input: $Code$: Code script
Output: $O$: Trajectory observation
$O\leftarrow\emptyset$
foreach  $action$ in $Code$  do
       $s_{\mathrm{last}}\leftarrow$ State Observation $(x,y,z,\theta)$
      Execute $action$
      if error executing $action$ then
             $O\leftarrow O$ + “Error in executing $action$ with error message $error$
       end if
      $s_{\mathrm{current}}\leftarrow$ State Observation $(x,y,z,\theta)$
      $\Delta s\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
      foreach $element$ in $\Delta s$ do
             if $element$ in orientation then
                   $O\leftarrow O$ + Rotate $element$ degrees in Yaw.
            else
                  if $element$ in NED axis then
                         $O\leftarrow O$ + “Move $element$ meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
                   end if
             end if
            $O\leftarrow O$ + “The UAV moves to $s_{\mathrm{current}}[x,y,z]$.”
       end foreach
end foreach
return $O$
Algorithm 2 State Transformation
IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the NL trajectory observation. The evaluator aims to identify deviations (if any) between the NL trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [40].
To improve the precision of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt. However, unlike the code generator, the evaluator does not produce code but instead generates evaluation feedback. To address this issue, we construct the evaluator’s system prompt with roles, rules, and references, as depicted in Fig. 3. Specifically, the roles define the LLM agent as an evaluator that compares the NL trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories. Additionally, the references are exemplar trajectories that serve as benchmarks to support the evaluation of NL trajectory observations. Furthermore, the system prompt requires the evaluation outcome to be informative and constructive by providing explanations about the depicted deviations. As demonstrated in Fig. 2, given a NL trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”). Consequently, the code generator is aware of the errors and could correctly refine the code in the subsequent refinement process.
Figure 3: The evaluation process begins by configuring the LLM agent (evaluator) via a system prompt framework that includes roles, rules, and references. Next, the trajectory observations and the task description are provided to the evaluator for comparison.
IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 2. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new NL trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs ”YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
V Experiment
The goal of this experiment is to validate the effectiveness of our proposed LLM-driven closed-loop framework in enhancing the reliability of LLM-driven UAV operation code generation. To achieve this, the experiment is designed with three primary goals: (1) to evaluate the performance of our closed-loop feedback and refinement framework in LLM-driven UAV tasks; (2) to analyze the impact of different design choices on the performance of our framework; (3) to assess the behavior of our framework under diverse task scenarios.
V-A Experimental Setup
V-A1 Environment
To ensure safe code execution without risking real-world failures, we conduct experiments in the “AirSim” simulator [41], which realistically simulates UAV (e.g., a quadcopter) dynamics and environmental interactions. During the experiment, UAV state information was fully accessible from the simulator in real time and utilized for closed-loop feedback and evaluation purposes.
V-A2 Tasks
We propose Basic and Advanced task sets to rigorously evaluate the performance of LLM-driven control code generation across varying levels of operational complexity [^1] The complete list of tasks is available at https://github.com/ai-uavsec/CLGSCE. The Basic set comprises 44 tasks adopted from GSCE [15]. Each task involves one or two intuitive discrete movements to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). In contrast, the Advanced set consists of 20 tasks that emulate real-world UAV deployment (e.g., “Take off and fly up 5 meters. You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left. Make sure the drone is oriented to the flying direction.”). These tasks demand reasoning about the objectives in the task description and require 6-19 UAV actions to reach the goal state. The advanced task set is intended to evaluate LLM’s multi-stage planning and complex spatial reasoning. Additionally, all tasks in both task sets were manually validated to avoid potential simulator-induced errors that could affect the evaluation process.
V-A3 Compared Methods
Three recently proposed methods [15, 25, 24] are selected as the baselines for comparison with our approach, including
  • GSCE: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement. [15]
  • LLM Self-feedback (Self-Refine): A single LLM performs both code generation and self-refinement. The LLM analyzes NL trajectory observations to detect deviations from the task and refines code based on its self-assessment.  [25]
  • Feedback with Numerical State Observations (Numerical): This method introduces a secondary LLM serving as evaluator designed according to section IV-D. The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement. [24]
All methods described above employ “o3-mini” (o3-mini-2025-01-31) [29] as the foundational LLM. The generated code was then deployed on a quadcopter in AirSim [41] to evaluate the performance of each method. To mitigate the randomness of LLM’s output [42], we repeat each task three times in the experiment.
V-B Evaluation Setup
V-B1 Metrics
The performance is evaluated using two metrics as GSCE [15]: Completeness and Success Rate (SR).
Completeness quantifies the proportion of goal states successfully reached, calculated as the ratio of the total count of correct actions to the total number of required actions of the ground truth trajectory in each task. The completeness offering insights into performance across the intermediate control process.
$Completeness=\frac{\sum\text{correct actions}}{\mathrm{len}(l)}$ (1)
SR is grounded in real-world operational requirements, wherein the UAV must fly the trajectory precisely as humans instructed. As formalized in Eq. 2, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory.
$Success=\begin{cases}1,&Completeness=1,\\ 0,&Completeness<1.\end{cases}$ (2)
V-B2 Ground Truth
The ground truth is represented by a list of state transitions. It is a vector of four elements: $[x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation. We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
V-C Result and Analysis
V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table I. For the Basic tasks, all methods showed remarkable performance. Specifically, the GSCE, Self-Refine, and Ours methods achieved $100\%$ in both SR and completeness, while the Numerical achieved $95.5\%$ SR and $97.3\%$ completeness. This demonstrates their ability to generate correct code for a few-stage planning and basic spatial reasoning. However, the slightly lower performance of the Numerical method suggests that relying on numerical state observations can occasionally lead to misinterpretation in evaluation, resulting in incomplete or incorrect refinements.
Performance on the Advanced tasks declines notably across all methods as a result of increased task complexity. The open-loop GSCE method achieves only $66.7\%$ SR, indicating that refinement becomes critical as task complexity increases. The Self-Refine method performs worse, with $50.0\%$ SR and $74.3\%$ completeness, proving that relying on a single LLM for both generation and self-evaluation amplifies bias [40], resulting in introducing additional errors in refinement. The Numerical method achieves $73.3\%$ SR and $92.4\%$ completeness, showing that it can detect some deviations; however, its numeric perspective still fails to provide reliable evaluation feedback for refinement. In contrast, our proposed method substantially outperforms all baselines, achieving $85.0\%$ SR and $98.5\%$ completeness on Advanced tasks. The overall results demonstrate that our proposed framework significantly enhances the reliability of LLM-driven UAV operation code generation, particularly for complex tasks that require multi-stage planning and complex spatial reasoning.
TABLE I: Average SR and Completeness of Proposed Methods
Base Task Advanced Task SR Completeness SR Completeness GSCE [15] $100.0\%$ $100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [25] $100.0\%$ $100.0\%$ $50.0\%$ $74.3\%$ Numerical [24] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $100.0\%$ $100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
V-C2 Failure Cases Analysis
We further analyze the failure cases (cases that $Success=0$) in the Advanced task set. The analysis focuses on three aspects: (1) the number of failures and average completeness for each failed task ID with the respective ground truth action length (Fig. 4); (2) the average completeness of each failed task ID (Fig. 5); (3) the amount of task that experienced failure and the average completeness of all failed tasks (Table II).
The results in Fig. 4 reveal a clear trend that all methods exhibit a notable increase in failures as task complexity increases (i.e., more likely to fail on tasks with longer action length). Specifically, the baseline GSCE method failed in $11$ out of $20$ tasks ($55\%$ of the Advanced set), with an average completeness of $77.4\%$. This indicates that the GSCE method can reason through a substantial portion of the task correctly, however, it still struggles to maintain full alignment with the objectives of the task. This result highlights the need for improvement in the completeness of failed tasks to reduce overall failure rates. The Self-Refine method exhibited the weakest performance, with failures on $17$ tasks ($85\%$ of Advanced tasks). Moreover, the failures in Self-Refine were the least complete among all methods, demonstrating that performing evaluation and refinement on a single LLM not only fails to correct errors but may also introduce more errors. The Numerical method failed on $9$ tasks ($45\%$ of Advanced tasks) and achieved an average completeness of $69.5\%$ only. These results indicate that numerical state observations are insufficient for reliable code refinement and may even lead to erroneous refinement.
Figure 4: Failure analysis across the four methods on the Advanced task set. The x-axis represents the task IDs that are sorted in ascending order based on the length of the task’s action steps. And the y-axis presents two metrics for each task: (1) the count of failures observed over three repetitions (left y-axis); (2) the length of the ground truth action steps (right y-axis).
Figure 5: Average completeness for each failed task on the Advanced task set over four methods. The smaller the value, the less complete the failed task is.
In contrast, our method achieves the best performance, failures occurred only on $4$ tasks ($20\%$ of Advanced tasks), and the completeness of failed tasks arose at $91.6\%$. Moreover, the completeness in Fig. 5 demonstrates that even when failures occur, our method produces code that remains closely aligned with task objectives. These findings underscore the effectiveness of our method in refining errors and improving the completeness of LLM-driven UAV operation, which contributes to an increased success rate in handling complex UAV operation tasks.
TABLE II: Amount of Task ID Experienced Failure and Average Completeness for All Failed Tasks
Task ID Amount $\downarrow$ Failed Tasks Completeness $\uparrow$
GSCE [15] $11$ $77.4\%$
Self-Refine [25] $17$ $48.5\%$
Numerical [24] $9$ $69.5\%$
Ours $\mathbf{4}$ $\mathbf{91.6\%}$
V-C3 Number of Refinement Iterations
[…]
Figure 6: Curves of average SR (left) and completeness (right) over the number of refinement iterations for the NL method on Advanced tasks.
V-D Evaluator Configuration Analysis
As the precision of the evaluation feedback directly influences the quality of code refinement, we also analyzed how different evaluator configurations affect evaluation precision. The evaluation precision is defined as the ratio of correct evaluations (i.e., correctly identifying whether the trajectory aligns with the task and providing a detailed explanation if deviations are identified) to the total number of evaluations conducted. To perform this analysis, we built a dataset from each Advanced task with one correct NL trajectory observation (Correct O) and one incorrect NL trajectory observation (Incorrect O). As presented in Table III, we tested four configurations of the evaluator system prompt: (1) roles; (2) roles $+$ rules; (3) roles $+$ references; (4) roles $+$ rules $+$ references. Moreover, each configuration was evaluated three times to reduce the randomness of LLM’s output [42].
TABLE III: Evaluator Configurations
Configuration (1) (2) (3) (4)
Roles
Rules